# Hilbert Space Fragmentation and Subspace Scar Time-Crystallinity in Driven Homogeneous Central-Spin Models

### v1

@@ B.2 XX type Hamiltonian with zero Zeeman energy
$\beta_{2p+1}^{m}$ $=(\beta_{2p}^{m}C_{m}-i2\gamma_{2p}^{m}S_{m}),$ (70)
@@ Appendix D Time-crystallinity in the XX central-spin model with coupling disorder
In this appendix, we study the robustness of the scar DTC behavior of the XX central-spin system without a Zeeman field in the presence of random coupling disorder. The disorder breaks the $\mathrm{I}_{t}^{2}$ symmetry of the Floquet operator, and therefore we cannot decompose the Hilbert space into different symmetry sectors. Therefore, we consider the full Hilbert space and the local spin-$z$ basis for the calculation of magnetization. We take a fully polarized satellite state ($\mathinner{|{J\downarrow}\rangle}=\mathinner{|{\uparrow_{1}\uparrow_{2}\uparrow_{3}\uparrow_{4}\uparrow_{5}}\rangle}\mathinner{|{\downarrow_{0})}\rangle}$ with 5 satellite spins and one central spin as the initial state. In this case, the average satellite spin magnetization is the same as the magnetization at each individual site. We sample the couplings from a Gaussian distribution with mean value $A_{xy}$ and variance (disorder strength) $\delta A_{xy}$. In Fig. 9 (a), we show the staggered average magnetization for the total satellite spin $\mathrm{I}_{t}^{z}$ over many Floquet cycles for different disorder coupling strengths. The average magnetization is about 0.5 in the absence of disorder (as expected from our study in the main text) or for a tiny amount of disorder across a large number of Floquet cycles. However, as the disorder strength is increased, the value of the average magnetization quickly goes to zero. Thus, disorder in the couplings adversely affects the robustness of time-crystalline behavior in the XX central-spin model.

### v2

@@ B.2 XX type Hamiltonian with zero Zeeman energy
$\beta_{2p+1}^{m}$ $=(\beta_{2p}^{m}C_{m}-i\gamma_{2p}^{m}S_{m}),$ (56)
@@ Appendix D Time-crystallinity in the XX central-spin model with coupling disorder
In this appendix, we study the robustness of the scar DTC behavior of the XX central-spin system without a Zeeman field in the presence of random coupling disorder. The disorder breaks the $\mathrm{I}_{t}^{2}$ symmetry of the Floquet operator, and therefore we cannot decompose the Hilbert space into different symmetry sectors. Therefore, we consider the full Hilbert space and the local spin-$z$ basis for the calculation of magnetization. We take a fully polarized satellite state ($\mathinner{|{J\downarrow}\rangle}=\mathinner{|{\uparrow_{1}\uparrow_{2}\uparrow_{3}\uparrow_{4}\uparrow_{5}}\rangle}\mathinner{|{\downarrow_{0})}\rangle}$ with 5 satellite spins and one central spin as the initial state. In this case, the average satellite spin magnetization is the same as the magnetization at each individual site. We sample the couplings from a Gaussian distribution with mean value $A_{xy}$ and variance (disorder strength) $\delta A_{xy}$.
@@
In Fig. 9 (a), we show the staggered average magnetization for the total satellite spin $\mathrm{I}_{t}^{z}$ over many Floquet cycles for different disorder coupling strengths. The average magnetization is about 0.5 in the absence of disorder (as expected from our study in the main text) or for a tiny amount of disorder across a large number of Floquet cycles. However, as the disorder strength is increased, the value of the average magnetization quickly goes to zero. Thus, disorder in the couplings adversely affects the robustness of time-crystalline behavior in the XX central-spin model.
@@
Appendix E Robustness of DTC in XXX and XXZ with magnetic field presence of coupling disorder
In this appendix, we explore the time-crystalline behavior of a fully polarized nuclear spin in the presence of disordered couplings for both the Heisenberg (XXX) and XXZ models, with a magnetic field applied to the central spin. The introduction of disorder renders the system non-integrable. The results obtained in this section can be interpreted as addressing two possible questions: (a) whether the time-crystalline behavior observed in the integrable system persists when the system becomes non-integrable, and (b) whether the time-crystalline behavior found under homogeneous (uniform) coupling persists in the presence of random inhomogeneities.
In Fig. 10 (a), (b), we show the staggered average magnetization for the total satellite spin $\mathrm{I}_{t}^{z}$ over many Floquet cycles for different coupling disorder strengths in the case of Heisenberg and XXZ interactions, respectively. The average magnetization is close to 0.5 in both cases. This shows that, for a fully polarized state, the time-crystalline behavior remains robust even in presence disorder coupling. In fact, this result is a special case of previous results in [54], where an MBL-DTC has been shown to exist in disordered driven XXZ and XXZ models.
Figure 10: Effect of random couplings in the XXX and XXZ central-spin systems with a Zeeman field of $B_{z}=300$ MHz on the central spin. The couplings are sampled from a Gaussian distribution with mean $A_{xyz}$ and variance (disorder strength) $\delta A_{xyz}$. (a) Staggered average satellite magnetization as a function of Floquet cycle for $A_{xyz}=1.3$ MHz for the XXX central-spin model. (b) Staggered average satellite magnetization as a function of Floquet cycle for $A_{xyz}=1.3$ MHz for the XXZ central-spin model, where $A_{z}=A_{xyz}\sqrt{0.9}$ and $A_{z}=A_{xyz}\sqrt{0.1}$. Here, $N=5$, the pulse error is $\theta=0.03\pi$, $\omega=1$ MHz, and the initial state is $\mathinner{|{J\downarrow}\rangle}$.